# Bayesian Posteriors with Stellar Population Synthesis on GPUs

### v1

@@ Acknowledgements
We thank the developers of NumPy (Van Der Walt et al., 2011), SciPy (Jones et al., 2001-2016), Jupyter (Ragan-Kelley et al., 2014), IPython (Pérez & Granger, 2007), JAX (Bradbury et al., 2018), conda-forge (conda-forge community, 2015), and Matplotlib (Hunter, 2007) for their extremely useful free software. While writing this paper we made extensive use of the Astrophysics Data Service (ADS) and arXiv preprint repository.
@@ Appendix A SPS model details
×$\qquad\times P_{\rm SFH}(\tau_{\rm age}^{i}|z,\theta_{\rm SFH})P_{\rm met}(Z^{j}|z,\theta_{\rm met})\;.$ (9)
@@ A.1 Star formation history
[…]
@@
where $M_{\star}^{\rm formed}$ is a normalization factor, $\alpha$ and $\beta$ are the two power-law indices describing the falling and rising slopes, respectively, while $\tau$ is related to the turnover time from one power-law to the other. We choose to use the double power-law due to its flexibility and because it has been shown to provide good fits to simulations (Pacifici et al., 2016; Diemer et al., 2017; Carnall et al., 2018; Mosleh et al., 2025). The parameters controlling the smooth SHF relation are then $\theta_{\rm SFR}^{\rm smooth}=(\log M_{\star}^{\rm formed},\tau,\alpha,\beta)$. The stellar age weights, $P_{\rm SFH}^{\rm smooth}$, are computed as the relative contribution to the total stellar mass $M_{\star}$ from stars in bin $i$ of stellar ages of range $\Delta\tau_{\rm age}^{i}$:
@@
×$\qquad\times\int_{t_{\rm lo}^{i}}^{t_{\rm hi}^{i}}dt^{\prime}\;\dot{M}_{\star}^{\rm smooth}(t^{\prime}|\theta_{\rm SFH})$ (12)
@@
In addition to the stellar population formed during “smooth” star formation, we separately model the contribution to the stellar population from a recent burst. A common approach to capturing the effect of bursty star formation is to use the power spectral density (PSD) to characterize short-timescale fluctuations (e.g., Caplar & Tacchella, 2019; Wang & Lilly, 2020a, b; Sun et al., 2025). In our SFH model, we instead formally decompose $P_{\rm SFH}(\tau)$ as follows:
@@
$\mathcal{T}(x)\equiv\left\{\begin{tabular}[]{ll}$(35/96)\left[1-(x/3)^{2}\right]^{3}$,&$\|x\|\leq 3$\\ $0$,&$\|x\|>3$\end{tabular}\right.\;.$ (16)
@@ A.2 Metallicity
The composite SED of a galaxy depends on the full metallicity distribution function (MDF), $P_{\rm SSP}(Z).$ However, galaxy photometry alone has limited constraining power on metallicity, and so it is common for SPS analyses to make the simplifying assumption that this PDF is a delta function centered at some $Z$ that is treated as a free parameter in the fit (e.g., Leja et al., 2020). In this work, we assume a lognormal form for the MDF, so that our model for metallicity includes two parameters, $\theta_{\rm met}=\{\log Z,\sigma_{\log Z}\}.$ The DSPS library includes alternative parametric forms for the MDF besides lognormal, including variations in which stellar age is correlated with metallicity, but variations on the assumed form of the MDF produce no appreciable change in the computational demands of the predictions, and so for purposes of this paper we limit our study to the lognormal form.
@@ A.3 Dust attenuation
To model the attenuation curve, we use the same functional form adopted in Salim et al. (2018):
@@
where $A_{\rm V}$ is the normalization defined at $\lambda_{\rm V}=5500\AA$, $k_{0}(\lambda)$ is the Calzetti function (Calzetti et al., 2000), and $D_{\lambda}$ is the standard Lorentzian-like Drude profile capturing the UV dust bump (Fitzpatrick & Massa, 1986).
@@
In Lower et al. (2022), the parameter $F_{\rm floor}$ was shown to improve the description of the effect of dust attenuation on mock SEDs in the Simba simulation (Davé et al., 2019), and in Hahn & Melchior (2025) it was demonstrated that $F_{\rm floor}$ likely plays an important role in mitigating biases in analyses of optical photometry.
@@ Appendix B Approximate photometry model
Figure B.2 shows the fractional differences between the apparent magnitude of a galaxy using the approximate versus the exact model. We have produced these model predictions using the gaussian transmission filters shown in Figure B.1. The range of resolution of these filters, defined as the ratio of the mean wavelength to the full width at half maximum (FWHM) of a transmission cure, $\lambda_{\rm mean}/{\rm FWHM}$, varies from $\sim 3.4$ for the left-most band to $\sim 25.5$ for the right-most one in Figure B.2. We perform the accuracy test for a large number of different galaxy properties $\theta_{\rm gal}$, by uniformly sampling from the parameter prior distributions (see Appendix C for more on the priors). The approximation is accurate to better than $0.1\%$ for a wide range of wavelengths. Note, however, that the level of accuracy of the approximation depends on the width of the transmission band: the broader the band, the less accurate this approximation will be. Thus the test we present here should be repeated before applying it to an analysis of a new dataset.
For the second test, we compare the posteriors from HMC using both the exact and the approximate photometry models when fitting an example galaxy with 12 model parameters. We perform this test using the exact same initial conditions for both fits and the results are shown in Figure B.3 for the full parameter space. We find that the posteriors from the approximate and exact photometry calculation are nearly indistinguishable. The approximation also speeds up the HMC calculation by greater than 30×30$\times$30 ×: for a converged chain with $3000$ steps, using the exact photometry takes 11 hours, while the runtime is 21 minutes when using the approximation.
Figure B.2: Fractional error between the approximate model for the apparent magnitude, $m_{\rm approx}$, and the exact model, $m_{\rm exact}$. Different color lines represent different choices of $\theta_{\rm gal}$ sampled uniformly from the priors.
@@ Appendix C Parameter bounding and prior distributions
In this appendix, we discuss how we construct the prior distributions in the Bayesian sampling algorithms we used in the main body of the text. Measurements of galaxy photometry provide a fundamentally limited amount of information about the fundamental properties of the galaxy, such as its star formation history (Ocvirk et al., 2006), and even in high signal-to-noise measurements the posterior information can be prior-dominated (Carnall et al., 2019; Leja et al., 2019a). For the purposes of this paper, we have opted for simple parametric families of PDFs for the prior, since our aims are primarily to assess matters of computational performance of running HMCs with stellar population synthesis on GPUs. As a much more flexible alternative, generative neural networks such as PopSED (Li et al., 2024) or pop-cosmos (Alsing et al., 2024) can capture complexity that is not possible with simple parametric families; depending on the size and complexity of the network, AI-based priors may come with a non-trivial computational cost and memory footprint when evaluating at scale that we do not consider here. Instead, we focus here on basic techniques for constructing families of priors while respecting custom bounds on the parameters of a physical model. First, we describe how we construct unbounded versions of the parameters in used in our fits. We then describe how we set up priors on the unbounded parameters, and finally transform the PDFs back into the bounded parameter space.
Consider an SPS parameter, $\theta,$ which on physical grounds should be bounded within some finite range. While running a gradient-based algorithm for either inference or optimization, there is no guarantee that a step in the direction of the gradient will respect the bounds. Some algorithms are naturally formulated to take such bounds into account (e.g., BFGS, Broyden, 1970; Fletcher, 1970; Goldfarb, 1970; Shanno, 1970), while others such as Adam (Kingma & Ba, 2014) are less adaptable. To address this issue, we construct unbounded parameters, $\tilde{\theta}$, which are related to $\theta$ through a bounding function, $\mathcal{F}_{\rm B},$ an invertible bijective transformation:
$\theta$ $=$ $\mathcal{F}_{\rm B}(\tilde{\theta})\;,$
@@
In Equation (19), the hyper-parameters of the sigmoid function for parameter $\theta$ are $p_{\theta}=\{y_{\rm lo}^{\theta},y_{\rm hi}^{\theta},k^{\theta},x_{0}^{\theta}\}$; these hyper-parameters need to be tuned in advance of running the sampler, and thereafter held fixed; the quantities $y_{\rm lo}^{\theta}$ and $y_{\rm hi}^{\theta}$ define the bounds on $\theta,$ while $k^{\theta}$ and $x_{0}^{\theta}$ should be tuned so that the behavior of $\mathcal{F}_{\rm B}$ is approximately linear in the range expected for the best-fit $\theta.$ More complex bounding functions than these can naturally arise when imposing constraints on a model; for example, the Diffstar parameterization for galaxy SFH (Alarcon et al., 2023) uses bounding functions that capture constraints on the parameter based on the physics of galaxy formation; bounding functions are typically lightweight, fast-evaluating functions that contribute negligibly to the computational demands of the Bayesian inference.
Once the bounding function $\mathcal{F}_{\rm B}$ is defined, it is straightforward to use it to construct families of priors in the space of unbounded parameters, $\tilde{\theta}$. We start with a simple assumption of approximately Gaussian priors on $\theta,$ draw $N$ samples, $\{{\theta}_{\rm i}\},$ and use $\mathcal{F}^{-1}_{\rm B}(\theta_{\rm i})$ to map into the unbounded space, $\{\tilde{\theta}_{\rm i}\}$. We parameterize $P(\tilde{\theta})$ using a skew-normal distribution:
@@ Appendix D Sampler details
D.1 Gradient-based sampler details
The two samplers we consider that utilize gradients are Hamiltonian Monte Carlo (HMC, Duane et al., 1987; Neal, 2011) and the No-U-Turn Sampler (NUTS, Hoffman & Gelman, 2014). HMC simulates a physical system with a potential and a kinetic term that explores the parameter space using the dynamics of the system. In particular, the leapfrog integrator is utilized by the algorithm for the system to move from one position to the next. The main benefit from using HMC over MCMC derives from the usage of gradients, which are incorporated into the kinetic term; having access to gradient information allows the sampler to construct trajectories through parameter space that are informed by the structure of the likelihood surface, reducing the noisy random-walk behavior of conventional MCMC samplers.
@@ D.1 Gradient-based sampler details
D.2 Sampler comparison
We have carried out a simple comparison of the results of each sampler to ensure that we derive comparable posteriors when adopting different sampling algorithms. We have run a fit to the same target data using both HMC and NUTS, as well as emcee (Foreman-Mackey et al., 2013), a widely used MCMC implementation. In Figure D.1 we present the parameter posteriors inferred from running HMC, NUTS and MCMC on the same target data. All chains have converged based on the $\hat{R}$ statistic values for all 12 model parameters. The runtimes for HMC, NUTS and MCMC are, respectively, $22$, $23$ and $29$ minutes. The MCMC sampler ran for $30,000$ steps with $32$ walkers, while HMC and NUTS ran for $3,000$ steps each, with additional $100$ steps of adaptation to tune the sampling parameters. The three samplers produce very similar contours. The contours of HMC and NUTS are visibly smoother than emcee due to the stochastic nature of MCMC sampling. The comparison to emcee gives non-trivial validation that our gradient-based posteriors are comparable to MCMC-based methods that are more widely used in the field of SPS; the comparison of HMC to NUTS also demonstrates that the HMC hyper-parameters that control the leapfrog integrator were tuned properly by the adaptation phase.
Figure D.1: Comparison between HMC, NUTS and MCMC, showing all 12 model parameters.

### v2

@@ Acknowledgements
We thank the developers of NumPy (numpy_ndarray), SciPy (scipy), Jupyter (jupyter), IPython (ipython), JAX (jax2018github), conda-forge (conda_forge_community_2015_4774216), and Matplotlib (matplotlib) for their extremely useful free software. While writing this paper we made extensive use of the Astrophysics Data Service (ADS) and arXiv preprint repository.
@@ Appendix A SPS model details
$\qquad\times P_{\rm SFH}(\tau_{\rm age}^{i}|z,\theta_{\rm SFH})P_{\rm met}(Z^{j}|z,\theta_{\rm met})\;.$ (9)
@@ A.1 Star formation history
$M_{\star}(z|\theta_{\rm gal})\equiv\int_{0}^{t_{\rm obs}}dt^{\prime}\;\dot{M}_{\star}(t^{\prime}|\theta_{\rm SFH})\;,$ (10)
@@
where $M_{\star}^{\rm formed}$ is a normalization factor, $\alpha$ and $\beta$ are the two power-law indices describing the falling and rising slopes, respectively, while $\tau$ is related to the turnover time from one power-law to the other. We choose to use the double power-law due to its flexibility and because it has been shown to provide good fits to simulations (Pacifici_2016; Diemer_2017; Carnall_2018; Mosleh2025). The parameters controlling the smooth SHF relation are then $\theta_{\rm SFR}^{\rm smooth}=(\log M_{\star}^{\rm formed},\tau,\alpha,\beta)$. The stellar age weights, $P_{\rm SFH}^{\rm smooth}$, are computed as the relative contribution to the total stellar mass $M_{\star}$ from stars in bin $i$ of stellar ages of range $\Delta\tau_{\rm age}^{i}$:
@@
$\qquad\times\int_{t_{\rm lo}^{i}}^{t_{\rm hi}^{i}}dt^{\prime}\;\dot{M}_{\star}^{\rm smooth}(t^{\prime}|\theta_{\rm SFH})$ (12)
@@
In addition to the stellar population formed during “smooth” star formation, we separately model the contribution to the stellar population from a recent burst. A common approach to capturing the effect of bursty star formation is to use the power spectral density (PSD) to characterize short-timescale fluctuations (e.g., Caplar2019_PSD; Wang2020_PSD_I; Wang2020_PSD_II; sun_etal24_bursty_sfh_psd_fire). In our SFH model, we instead formally decompose $P_{\rm SFH}(\tau)$ as follows:
@@
[…]
@@ A.2 Metallicity
The composite SED of a galaxy depends on the full metallicity distribution function (MDF), $P_{\rm SSP}(Z).$ However, galaxy photometry alone has limited constraining power on metallicity, and so it is common for SPS analyses to make the simplifying assumption that this PDF is a delta function centered at some $Z$ that is treated as a free parameter in the fit (e.g., leja_etal_2020_smf_census). In this work, we assume a lognormal form for the MDF, so that our model for metallicity includes two parameters, $\theta_{\rm met}=\{\log Z,\sigma_{\log Z}\}.$ The DSPS library includes alternative parametric forms for the MDF besides lognormal, including variations in which stellar age is correlated with metallicity, but variations on the assumed form of the MDF produce no appreciable change in the computational demands of the predictions, and so for purposes of this paper we limit our study to the lognormal form.
@@ A.3 Dust attenuation
To model the attenuation curve, we use the same functional form adopted in Salim2018Dust:
@@
where $A_{\rm V}$ is the normalization defined at $\lambda_{\rm V}=5500\AA$ , $k_{0}(\lambda)$ is the Calzetti function (calzetti00), and $D_{\lambda}$ is the standard Lorentzian-like Drude profile capturing the UV dust bump (fitzpatrick_massa_1986).
@@
In Lower2022Dust, the parameter $F_{\rm floor}$ was shown to improve the description of the effect of dust attenuation on mock SEDs in the Simba simulation (dave_etal19_simba), and in hahn_melchior_2025_clumpy_dust it was demonstrated that $F_{\rm floor}$ likely plays an important role in mitigating biases in analyses of optical photometry.
@@ Appendix B Approximate photometry model
Figure B.2 shows the fractional differences between the apparent magnitude of a galaxy using the approximate versus the exact model. We have produced these model predictions using the gaussian transmission filters shown in Figure B.1. The range of resolution of these filters, defined as the ratio of the mean wavelength to the full width at half maximum (FWHM) of a transmission cure, $\lambda_{\rm mean}/{\rm FWHM}$, varies from $\sim 3.4$ for the left-most band to $\sim 25.5$ for the right-most one in Figure B.2. We perform the accuracy test for a large number of different galaxy properties $\theta_{\rm gal}$, by uniformly sampling from the parameter prior distributions (see Appendix C for more on the priors). The approximation is accurate to better than $0.06$ magnitude for a wide range of wavelengths. Note, however, that the level of accuracy of the approximation depends on the width of the transmission band: the broader the band, the less accurate this approximation will be. Thus the test we present here should be repeated before applying it to an analysis of a new dataset. As a test case for the applicability of our approximate model to more realistic transmission bands, we performed a test that is similar to the above using the 6 LSST bands $(u,g,r,i,z,y)$. We computed theoretical photometry predictions for various sets of model parameters comparing the exact with the approximate photometry model, in both cases using the LSST bands, and the results are presented in Figure B.3. As we can see, the approximate model works well in this case as well, always resulting in magnitude differences less than $0.03$.
For the second test, we compare the posteriors from HMC using both the exact and the approximate photometry models when fitting an example galaxy with 12 model parameters. We perform this test using the exact same initial conditions for both fits and the results are shown in Figure B.4 for the full parameter space. We find that the posteriors from the approximate and exact photometry calculation are nearly indistinguishable. The approximation also speeds up the HMC calculation by greater than $30\times$: for a converged chain with $3000$ steps, using the exact photometry takes 11 hours, while the runtime is 21 minutes when using the approximation.
Figure B.2: Difference between the approximate model for the apparent magnitude, $m_{\rm approx}$, and the exact model, $m_{\rm exact}$. Different color lines represent different choices of $\theta_{\rm gal}$ sampled uniformly from the priors.
Figure B.3: Similar to Figure B.2 but this time using the 6 LSST transmission bands $(u,g,r,i,z,y)$ for a more realistic test on the accuracy of the approximate photometry compared to the exact model.
@@ Appendix C Parameter bounding and prior distributions
In this appendix, we discuss how we construct the prior distributions in the Bayesian sampling algorithms we used in the main body of the text. Measurements of galaxy photometry provide a fundamentally limited amount of information about the fundamental properties of the galaxy, such as its star formation history (ocvirk_etal06), and even in high signal-to-noise measurements the posterior information can be prior-dominated (Carnall_2019_SFH; leja_etal19_how_to_measure2). For the purposes of this paper, we have opted for simple parametric families of PDFs for the prior, since our aims are primarily to assess matters of computational performance of running HMCs with stellar population synthesis on GPUs. As a much more flexible alternative, generative neural networks such as PopSED (li_etal24_popsed) or pop-cosmos (alsing_etal24_pop_cosmos) can capture complexity that is not possible with simple parametric families; depending on the size and complexity of the network, AI-based priors may come with a non-trivial computational cost and memory footprint when evaluating at scale that we do not consider here. Instead, we focus here on basic techniques for constructing families of priors while respecting custom bounds on the parameters of a physical model. First, we describe how we construct unbounded versions of the parameters used in our fits. We then describe how we set up priors on the unbounded parameters, and finally transform the PDFs back into the bounded parameter space.
Consider an SPS parameter, $\theta,$ which on physical grounds should be bounded within some finite range. While running a gradient-based algorithm for either inference or optimization, there is no guarantee that a step in the direction of the gradient will respect the bounds. Some algorithms are naturally formulated to take such bounds into account (e.g., BFGS, broyden_1970_B_in_BFGS; fletcher_1970_F_in_BFGS; goldfarb_1970_G_in_BFGS; shanno_1970_S_in_BFGS), while others such as Adam (kingma_ba_adam_2015) are less adaptable. To address this issue, we construct unbounded parameters, $\tilde{\theta}$, which are related to $\theta$ through a bounding function, $\mathcal{F}_{\rm B},$ an invertible bijective transformation:
[…]
@@
In Equation (19), the hyper-parameters of the sigmoid function for parameter $\theta$ are $p_{\theta}=\{y_{\rm lo}^{\theta},y_{\rm hi}^{\theta},k^{\theta},x_{0}^{\theta}\}$; these hyper-parameters need to be tuned in advance of running the sampler, and thereafter held fixed; the quantities $y_{\rm lo}^{\theta}$ and $y_{\rm hi}^{\theta}$ define the bounds on $\theta,$ while $k^{\theta}$ and $x_{0}^{\theta}$ should be tuned so that the behavior of $\mathcal{F}_{\rm B}$ is approximately linear in the range expected for the best-fit $\theta.$ More complex bounding functions than these can naturally arise when imposing constraints on a model; for example, the Diffstar parameterization for galaxy SFH (diffstar_alarcon_etal23) uses bounding functions that capture constraints on the parameter based on the physics of galaxy formation; bounding functions are typically lightweight, fast-evaluating functions that contribute negligibly to the computational demands of the Bayesian inference.
[…]
@@ Appendix D Sampler details
D.1 Simulated galaxy sample
The galaxy sample we use in this work to fit our model to is simulated. In particular, we use the SPS model that is described in Appendix A to both produce the simulated dataset and to make model predictions while running our inference pipeline. The simulated sample consists of 20000 galaxies, each of which is described by a set of 12 parameters, sampled from the prior distributions, to characterize the galaxy’s star-formation history, mass-metallicity relation, dust attenuation, and burstiness properties. In correspondence with these parameter sets, we generate mock photometry for each galaxy using the mathematical framework we describe in Section 2 and Appendix A with the transmission bands presented in Figure B.1. Thus, each galaxy is also described by a 20-point photometry vector. In addition, we design a diagonal gaussian covariance matrix to accompany this data vector, defined as $\mathcal{C}=0.2\times\mathbb{I}_{20}$, where $\mathbb{I}_{20}$ is the $20\times 20$ identity matrix. The gaussian width $\sigma_{\rm phot}=0.2$ of this covariance is chosen as a simple but reasonable value of the uncertainty on the photometry of each galaxy.
D.2 Gradient-based sampler details
The two samplers we consider that utilize gradients are Hamiltonian Monte Carlo (HMC, duan_etal87_hmc; neal_2011_hmc_chapter) and the No-U-Turn Sampler (NUTS, hoffman_gelman_2014_nuts). HMC simulates a physical system with a potential and a kinetic term that explores the parameter space using the dynamics of the system. In particular, the leapfrog integrator is utilized by the algorithm for the system to move from one position to the next. The main benefit from using HMC over MCMC derives from the usage of gradients, which are incorporated into the kinetic term; having access to gradient information allows the sampler to construct trajectories through parameter space that are informed by the structure of the likelihood surface, reducing the noisy random-walk behavior of conventional MCMC samplers.
@@ D.2 Gradient-based sampler details
D.3 Sampler comparison
We have carried out a simple comparison of the results of each sampler to ensure that we derive comparable posteriors when adopting different sampling algorithms. We have run a fit to the same target data using both HMC and NUTS, as well as emcee (EMCEEpaper), a widely used MCMC implementation. In Figure D.1 we present the parameter posteriors inferred from running HMC, NUTS and MCMC on the same target data. All chains have converged based on the $\hat{R}$ statistic values for all 12 model parameters. The runtimes for HMC, NUTS and MCMC are, respectively, $22$, $23$ and $29$ minutes. The MCMC sampler ran for $30,000$ steps with $32$ walkers, while HMC and NUTS ran for $3,000$ steps each, with additional $100$ steps of adaptation to tune the sampling parameters. The three samplers produce very similar contours. The contours of HMC and NUTS are visibly smoother than emcee due to the stochastic nature of MCMC sampling. The comparison to emcee gives non-trivial validation that our gradient-based posteriors are comparable to MCMC-based methods that are more widely used in the field of SPS; the comparison of HMC to NUTS also demonstrates that the HMC hyper-parameters that control the leapfrog integrator were tuned properly by the adaptation phase.
Figure D.1: Comparison between HMC, NUTS and $MC$, showing all 12 model parameters.